### v1

@@ 1. Introduction
Building 3D worlds and objects for Virtual Reality (VR) has become increasingly accessible, enabling users to immerse themselves in detailed environments and also actively (co-)create them. Recent advances have also seen the integration of generative Artifical Intelligence (AI) in this process (De La Torre et al., 2024), leveraging text prompts and other inputs to generate elaborate 3D objects and even entire 3D worlds. However, these emerging approaches often rely on a ”set-it-and-forget-it” approach, where users input a prompt and receive a fully realized 3D object or world with limited opportunities for customization or interaction. While this trend toward AI-driven generation can democratize access to high-quality 3D objects, it challenges human agency and shifts control from humans to AI.
@@
To empower users to get more involved in the design process, research has started exploring co-creative systems, which enable users to collaborate with AI in a creative process (Davis, 2021). These systems have been applied across a variety of domains, including music, storytelling, game design, and visual arts (Liapis et al., 2013; Lucas and Martinho, 2017; Davis et al., 2016; Oh et al., 2018; Lin et al., 2020). In recent years, there has been growing interest in extending such co-creative systems to 3D object co-creation. AI systems can now generate and modify 3D objects based on user inputs such as text prompts or sketches (Poole et al., 2022; Lin et al., 2023; Michel et al., 2021; Jun and Nichol, 2023; Mikaeili et al., 2023; Lun et al., 2017), and some systems even allow for interactive refinement of AI outputs (Liu et al., 2018; Urban Davis et al., 2021).
@@
Despite these advances, a critical question remains largely unexplored: how should the actions of a co-creative tool for object generation be represented in VR? This question touches on fundamental aspects of co-creative interactions and raises several essential considerations: (1) Human collaborative work often involves asynchronous or gradual changes, making immediate modifications potentially counterintuitive. Should AI-driven changes to 3D objects be presented incrementally, mirroring the rhythm of human collaboration, or is immediate modification more appropriate in the context of co-creation? (2) Drawing from literature on non-player characters (NPCs), embodiment is often deemed crucial for effective interaction. However, does this hold true for a co-creative system focused on object generation, which may not inherently require an embodied visualization as NPCs do? Does the user perception of the system change from a tool to a collaborator only through embodiment, and if yes, what are potential downsides? (3) Decades of HCI research and design practices emphasize the importance of highlighting focus or activity (e.g., showing all cursors in collaborative text editing tools or speaker indicators in video conferencing). Is visually indicating where and how AI modifications occur sufficient for co-creative object generation systems, or do these systems necessitate new ways of co-creative interaction?
@@
Addressing these questions is pivotal to advancing the design of co-creative systems for 3D object and world generation. This paper hence contributes to answering these questions by investigating different representations of AI actions in the context of co-creating 3D objects in VR (see Figure 1). Specifically, we examine the aforementioned three aspects of co-creating a 3D object in VR in a Wizard-of-Oz user study on (1) perception of embodiment, (2) incremental visualization of co-creative contributions, and (3) effect of highlighting the area where the AI is going to perform a modification. Our findings show that co-creating with an embodied AI significantly influences the perceived AI contribution to the created model. Further, highlighting does not increase predictability or communication with the AI system but decreases users’ enjoyment and a perceived partnership with the AI. Finally, users pay more attention to the AI when it uses an incremental generation.
@@ 2. Related Work
In the following, we discuss related work in the areas of 3D modeling in VR, AI for 3D object generation, co-creative systems in general, embodied AI, and AI contribution visualization.
@@
Three-dimensional modeling has been a topic of interest in computer research since as early as the 1970s (Clark, 1976), with HoloSketch (Deering, 1995) being among the earliest contributions to 3D modeling software for immersive environments. In general, there are several ways to approach 3D design in VR: For instance, sketching, though intuitive and quick, suffers from a lack of haptic feedback mid-air. To mitigate inaccuracies, some systems smooth out hand-drawn lines in post-processing (Wesche and Seidel, 2001), feature the use of real physical objects in AR as guides (Reipschläger and Dachselt, 2019), use haptic feedback of the controller as assistance (Elsayed et al., 2020) or let the user sketch on 2D surfaces and then transform the lines in 3D space (Arora et al., 2018; Jackson and Keefe, 2016; Wesche and Seidel, 2001; Dorta et al., 2016). Similar problems arise with digital sculpting, in which the user shapes desired forms out of primitives and pre-included models using virtual sculpting tools (Sketch, 2020). Other approaches include mesh modeling (Reipschläger and Dachselt, 2019; Butterworth et al., 1992) and assembly from primitives (Deering, 1995; Reipschläger and Dachselt, 2019; Do and Lee, 2010). Additionally, there also exist several commercially available tools for 3D design in VR that allow users to paint 3D artworks (LLC, 2016), design objects by free-hand drawing (Sketch, 2020), and modify the mesh directly (Toff, 2017). Closely related to this paper, GetWild (Wong et al., 2022) is a VR environment-editing system incorporating AI-generated models into the creation process to speed up the design pipeline and increase accessibility for users with little modeling experience.
@@ 2.2. AI for 3D Object Generation and Manipulation
While AI research for 2D image synthesis has made notable progress in the past few years, the field of 3D object generation has faced more difficulties due to limitations of the different representation modes of 3D data (Shi et al., 2023; Li et al., 2023) as well as a lack of text-to-3D training data for deep learning models (Li et al., 2023). Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) are a recently developed approach to representing 3D data, originally serving the purpose of 3D image reconstruction. Due to their ability to bridge the gap between 2D and 3D data, NeRFs can be combined with powerful text-to-image generative models to construct advanced text-to-3D generative models (Li et al., 2023). This has led to a recent upsurge in AI tools for 3D object generation, such as DreamFusion (Poole et al., 2022) and Magic3D (Lin et al., 2023). However, outputs of NeRF-based 3D generation are difficult to use in downstream 3D applications since a lot of 3D graphics software requires standard data representations like meshes (Li et al., 2023). Shap-E (Jun and Nichol, 2023) is a recent work demonstrating object generation based on a text prompt, with a mesh output being created within less than two minutes.
Though generative models like the ones mentioned above are already rather useful for exploring 3D design ideas, it can be difficult for users to formulate a textual prompt that will produce an output meeting all of the user’s envisioned criteria. This one-step approach often forces the user to make the AI generate an entirely new object based on a rephrased prompt. For this reason, efforts have been made to develop AI tools that can modify a given 3D object according to a user’s instructions.
SKED (Mikaeili et al., 2023) uses sketch-based instructions to enable interactive editing of 3D objects, whereas ShapeCrafter (Fu et al., 2023) generates a 3D object from a text-based description and lets the user refine object characteristics by recursively adding descriptive phrases. Similarly, Text2Mesh (Michel et al., 2021) can change the style of 3D objects: Given an input mesh and a text prompt, it modifies color and geometry details that adhere to the user-specified style. Additionally, SPAGHETTI (Hertz et al., 2022) allows users to select parts of a 3D object and perform rough local transformations on them, based on which the generative model creates an edited version of the same object.
Although previous works include interactive elements similar to co-creative systems, none leverage the intuitive and expressive potential of spatial 3D interaction and presentation offered by VR. Instead, these systems rely on indirect instructions via text or 2D interface interactions. In contrast, this work specifically focuses on VR environments.
@@ 2.3. Co-creative Systems
A stream of research has also started to investigate co-creative systems for editing 3D objects. For instance, Liu et al. (Liu et al., 2018) have proposed a system in which users can take turns creating and editing 3D objects with an AI. Closely related, Calliope (Urban Davis et al., 2021) has explored interaction possibilities of using generative adversarial networks as an active collaborator in VR. However, it did not investigate the specific influences of different representation modes of such a collaborator on user perception. Other than Calliope, co-creative systems have typically not been studied in VR environments but rather in physical or digital 2D ones. We want to explore whether findings from those areas also apply to co-creative systems in VR.
@@ 2.4. Embodied AI
Embodiment can be defined in several ways, though the most basic form of an embodied system has been conceptualized by Ziemke (Ziemke, 2003) as a system that can act on and be acted on by its environment. Other notions summarized by Ziemke include that of physical embodiment, which necessitates a physical body, and organismoid embodiment, which further constrains the physical body to resemble that of a living being. Guckelsberger et al. (Guckelsberger et al., 2021) extend Ziemke’s typology with the concept of virtual embodiment, which requires a virtual body that can act on and be acted on by a virtual environment.
Research has shown that the virtual embodiment of agents can improve motivation, positive attitude, and collaborative experiences in creative systems (Baylor, 2009; Rezwana and Maher, 2022b; Davis et al., 2016). Additionally, organismoid embodiment enhances identification, empathy, and perceived creativity, and is considered essential for co-creativity due to the unique perspective it provides (Guckelsberger et al., 2021). Also Kim et al. (Kim et al., 2018) investigated the effect of embodiment on the perception of a digital AR assistant. Moruzzi (Moruzzi, 2022) found that AI embodiment increased perceived creativity in a collaborative artistic process, while Rezwana and Maher (Rezwana and Maher, 2023) showed that AI personification influenced users’ perceptions of the AI as an independent collaborator and affected their ethical views on ownership of co-created artifacts.
Studies on the effects of embodiment in VR have mainly examined AI agents without co-creativity, while studies on the embodiment of co-creative AI have mainly been carried out in non-VR environments. We combine both promising directions of previous works by studying embodiment in the context of a co-creative object-generating agent embedded in VR.
@@ 2.5. Non-player Characters in Games
In the domain of game development, the design of autonomous human-like agents, so called Non-Player Characters (NPCs), was addressed in the context of believability (Bates, 1994). Recent research explores the potential of NPCs in virtual reality (Yin and Xiao, 2024) and of LLM-based agents with more human-like reasoning, planning, and execution abilities (Park et al., 2023). Moreover, in a systematic literature review, Wittmann and Morschheuser (2022) discuss NPC design patterns and how to transfer those to the design of AI systems for human-AI collaboration. The authors point out six relevant focus fields namely the NPC responsiveness, appearance, communication patterns, emotional aspects, behavioral characteristics, and player-NPC and NPC-NPC team structures.
Unlike NPCs, which are inherently embodied as avatars, co-creative tools for object building do not inherently require such embodiment. This raises important questions about the role of embodiment in such tools and its interaction with factors like the highlighting and visualization of changes, which are the focus of this paper.
@@ 2.6. AI Contribution Visualization
Thus far, no empirical studies have investigated the effects of such an intentionally incremental visualization of AI contributions, and especially not separately to effects of AI embodiment. We therefore want to address this lack of research by conducting a study in which both embodiment as well as perceptual evidence in the form of incremental contribution visualization are examined in the context of a co-creative VR system.
@@ 3. Methodology
How does highlighting affect the co-creative collaboration experience, in particular regarding the predictability of AI? Based on decades of HCI research and design practices that emphasize the importance of highlighting focus or activity, we hypothesize that highlighting will have a positive effect on the perceived communication quality, as it is a well-known mechanism to guide users’ attention. For the same reason, we also expect that with highlighting, users will pay more attention to the AI, and the clarity of the AI intentions will be perceived better.
@@
How does the embodiment of an AI affect the co-creative collaboration experience, in particular regarding the perceived supportiveness, efficiency, and competence of the AI? Research on NPCs and games highlights embodiment as a critical factor in interaction, making embodiment a potentionally promising factor also for co-creative tools in VR. We thus expect that embodiment will affect the perceived supportiveness, efficiency, and competence of the AI, the perceived value and creativity of AI outputs, measures concerning collaborative experience, appeal, and behavioral engagement, as well as the previously mentioned contribution attribution and users’ perceived closeness to the AI.
Showing the different IVs with an example of a chair. Left are four images illustrating the difference for disembodied AI. Right are four images showing the VR scene with an embodied AI.
@@ 3.1. Task
At the beginning of a modeling session, a 3D object appears on the podium and the timer starts running. As long as the timer was running, the user could modify the object by either painting its texture (see Figure 2), add elements to the existing object (see Figure 2), or instructing the AI to modify (see Figure 2). We chose painting as the users’ main design responsibility as it allowed for a consistent engagement with the 3D object during the AI-driven modifications and ensured comparable levels of object alteration and creative expression while avoiding interference. For the modifications, users were always controlling if and when any modifications to the object should be executed (including being able to revert the AI’s modification). We simulated the AI agent in a Wizard-of-Oz approach, in the sense that its contributions to the 3D objects were not the results of an actual generative process but instead object additions and replacements created by human 3D artists in advance.
Further, the creative contributions of the AI agent are not designed to align with the contributions of the human partner, but instead emerge completely independently of the human’s behavior. This contribution style is typical for what Rezwana and Maher (Rezwana and Maher, 2022a) refer to as provoking agents (as opposed to pleasing agents), which evoke creativity in their human partners by providing divergent ideas and alternative explorations. While the AI was active, the user could not give another modification instruction to the AI.

### v3

@@ 1. Introduction
Building 3D worlds and objects for Virtual Reality (VR) has become increasingly accessible, enabling users to immerse themselves in detailed environments and actively (co-)create them. Recent advances have also seen the integration of generative Artificial Intelligence (AI) in this process (De La Torre et al., 2024), leveraging text prompts and other inputs to generate elaborate 3D objects and even entire 3D worlds. However, these emerging approaches often rely on a ”set-it-and-forget-it” approach, where users input a prompt and receive a fully realized 3D object or world with limited opportunities for customization or interaction. While this trend toward AI-driven generation can democratize access to high-quality 3D objects, it challenges human agency and shifts control from humans to AI.
@@
To empower users to get more involved in the design process, a stream of research has started exploring co-creative systems, which enable users to collaborate with AI in a creative process (Davis, 2021). These systems have been applied across a variety of domains, including music, storytelling, game design, and visual arts (Liapis et al., 2013; Lucas and Martinho, 2017; Davis et al., 2016; Oh et al., 2018; Lin et al., 2020). There has been growing interest in extending such co-creative systems to 3D object co-creation in recent years. AI systems can now generate and modify 3D objects based on user inputs such as text prompts or sketches (Poole et al., 2022; Lin et al., 2023; Michel et al., 2021; Jun and Nichol, 2023; Mikaeili et al., 2023; Lun et al., 2017), and some systems even allow for interactive refinement of AI outputs (Liu et al., 2018; Urban Davis et al., 2021).
@@
Despite these advances, a critical question remains largely unexplored: how should the actions of a co-creative tool for object generation be represented in VR? This question touches on fundamental aspects of co-creative interactions and raises several essential considerations: (1) Human collaborative work often involves asynchronous or gradual changes, making immediate modifications potentially counterintuitive. Should AI-driven changes to 3D objects be presented incrementally, mirroring the rhythm of human collaboration, or is immediate modification more appropriate in co-creation? (2) Drawing from the literature on non-player characters (NPCs), the embodiment is often deemed crucial for effective interaction. However, does this hold for a co-creative system focused on object generation, which may not inherently require an embodied visualization as NPCs do? Does the user perception of the system change from a tool to a collaborator only through embodiment, and if yes, what are the potential downsides? (3) Decades of HCI research and design practices emphasize highlighting focus or activity (e.g., showing all cursors in collaborative text editing tools or speaker indicators in video conferencing). Is visually indicating where and how AI modifications occur sufficient for co-creative object generation systems, or do these systems necessitate new ways of co-creative interaction?
@@
Addressing these questions is pivotal to advancing the design of co-creative systems for 3D objects and world generation. This paper contributes to answering these questions by investigating different representations of AI actions in the context of co-creating 3D objects in VR (see Figure 1). Specifically, we examine the aforementioned three aspects of co-creating a 3D object in VR in a Wizard-of-Oz user study on (1) perception of embodiment, (2) incremental visualization of co-creative contributions, and (3) effect of highlighting the area where the AI is going to perform a modification. Our findings show that co-creating with an embodied AI significantly influences the perceived AI contribution to the created model. Further, highlighting does not increase predictability or communication with the AI system but decreases users’ enjoyment and a perceived partnership with the AI. Finally, users pay more attention to the AI when it uses an incremental generation.
@@ 2. Related Work
In the following, we discuss related work in 3D modeling in VR, AI for 3D object generation, co-creative systems in general, embodied AI, and AI contribution visualization.
@@
Three-dimensional modeling has been a topic of interest in computer research since as early as the 1970s (Clark, 1976), with HoloSketch (Deering, 1995) being among the earliest contributions to 3D modeling software for immersive environments. In general, there are several ways to approach 3D design in VR. For instance, sketching, though intuitive and quick, suffers from a lack of haptic feedback mid-air. Some systems mitigate inaccuracies by smoothing out hand-drawn lines in post-processing (Wesche and Seidel, 2001), featuring the use of real physical objects in AR as guides (Reipschläger and Dachselt, 2019), using haptic feedback of the controller as assistance (Elsayed et al., 2020) or letting the user sketch on 2D surfaces and then transform the lines in 3D space (Arora et al., 2018; Jackson and Keefe, 2016; Wesche and Seidel, 2001; Dorta et al., 2016). Similar problems arise with digital sculpting, in which the user shapes desired forms out of primitives and pre-included models using virtual sculpting tools (Sketch, 2020). Other approaches include mesh modeling (Reipschläger and Dachselt, 2019; Butterworth et al., 1992) and assembly from primitives (Deering, 1995; Reipschläger and Dachselt, 2019; Do and Lee, 2010). Additionally, there also exist several commercially available tools for 3D design in VR that allow users to paint 3D artworks (LLC, 2016), design objects by free-hand drawing (Sketch, 2020), and modify the mesh directly (Toff, 2017). Closely related to this paper, GetWild (Wong et al., 2022) is a VR environment-editing system incorporating AI-generated models into the creation process to speed up the design pipeline and increase accessibility for users with little modeling experience.
@@ 2.2. AI for 3D Object Generation and Manipulation
While AI research for 2D image synthesis has made notable progress in the past few years, the field of 3D object generation has faced more difficulties due to limitations of the different representation modes of 3D data (Shi et al., 2023; Li et al., 2023) as well as a lack of text-to-3D training data for deep learning models (Li et al., 2023). Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) is a recently developed approach to representing 3D data, initially serving the purpose of 3D image reconstruction. Due to their ability to bridge the gap between 2D and 3D data, NeRFs can be combined with powerful text-to-image generative models to construct advanced text-to-3D generative models (Li et al., 2023). This has led to a recent upsurge in AI tools for 3D object generation, such as DreamFusion (Poole et al., 2022) and Magic3D (Lin et al., 2023). However, outputs of NeRF-based 3D generation are difficult to use in downstream 3D applications since a lot of 3D graphics software requires standard data representations like meshes (Li et al., 2023). Shap-E (Jun and Nichol, 2023) is a recent work demonstrating object generation based on a text prompt, with a mesh output created within less than two minutes.
Though generative models like the ones mentioned above are already helpful for exploring 3D design ideas, it can be difficult for users to formulate a textual prompt that will produce an output meeting all of the user’s envisioned criteria. This one-step approach often forces the user to make the AI generate an entirely new object based on a rephrased prompt. For this reason, efforts have been made to develop AI tools that can modify a given 3D object according to a user’s instructions.
SKED (Mikaeili et al., 2023) uses sketch-based instructions to enable interactive editing of 3D objects. In contrast, ShapeCrafter (Fu et al., 2023) generates a 3D object from a text-based description and lets the user refine object characteristics by recursively adding descriptive phrases. Similarly, Text2Mesh (Michel et al., 2021) can change the style of 3D objects: Given an input mesh and a text prompt, it modifies color and geometry details that adhere to the user-specified style. Additionally, SPAGHETTI (Hertz et al., 2022) allows users to select parts of a 3D object and perform rough local transformations on them, based on which the generative model creates an edited version of the same object.
Although previous works include interactive elements similar to co-creative systems, none leverage the intuitive and expressive potential of spatial 3D interaction and presentation offered by VR. Instead, these systems rely on indirect instructions via text or 2D interface interactions. In contrast, this work focuses explicitly on VR environments.
@@ 2.3. Co-creative Systems
A stream of research has also started to investigate co-creative systems for editing 3D objects. For instance, Liu et al. (Liu et al., 2018) have proposed a system where users can take turns creating and editing 3D objects with an AI. Closely related, Calliope (Urban Davis et al., 2021) has explored interaction possibilities of using generative adversarial networks as an active collaborator in VR. However, it did not investigate the specific influences of different representation modes of such a collaborator on user perception. Other than Calliope, co-creative systems are typically not studied in VR environments but in physical or digital 2D ones. We want to explore whether findings from those areas also apply to co-creative systems in VR.
@@ 2.4. Embodied AI
Embodiment can be defined in several ways, though Ziemke (Ziemke, 2003) has conceptualized the most basic form of an embodied system as a system that can act on and be acted on by its environment. Other notions include physical embodiment, which necessitates a physical body, and organismoid embodiment, which further constrains the physical body to resemble a living being. Guckelsberger et al. (Guckelsberger et al., 2021) extend Ziemke’s typology with the concept of virtual embodiment, which requires a virtual body that can act on and be acted on by a virtual environment.
Research has shown that the virtual embodiment of agents can improve motivation, positive attitude, and collaborative experiences in creative systems (Baylor, 2009; Rezwana and Maher, 2022b; Davis et al., 2016). Additionally, organismoid embodiment enhances identification, empathy, and perceived creativity and is considered essential for co-creativity due to the unique perspective it provides (Guckelsberger et al., 2021). Also, Kim et al. (Kim et al., 2018) investigated the effect of embodiment on the perception of a digital AR assistant. Moruzzi (Moruzzi, 2022) found that AI embodiment increased perceived creativity in a collaborative artistic process, while Rezwana and Maher (Rezwana and Maher, 2023) showed that AI personification influenced users’ perceptions of the AI as an independent collaborator and affected their ethical views on ownership of co-created artifacts.
Studies on the effects of embodiment in VR have mainly examined AI agents without co-creativity. In contrast, studies on the embodiment of co-creative AI have mainly been carried out in non-VR environments. We combine both promising directions of previous works by studying embodiment in the context of a co-creative object-generating agent embedded in VR.
@@ 2.5. Non-player Characters in Games
In the domain of game development, the design of autonomous human-like agents, so-called Non-Player Characters (NPCs), was addressed in the context of believability (Bates, 1994). Recent research explores the potential of NPCs in virtual reality (Yin and Xiao, 2024) and of LLM-based agents with more human-like reasoning, planning, and execution abilities (Park et al., 2023). Moreover, in a systematic literature review, Wittmann and Morschheuser (2022) discuss NPC design patterns and how to transfer those to the design of AI systems for human-AI collaboration. The authors point out six relevant focus fields: NPC responsiveness, appearance, communication patterns, emotional aspects, behavioral characteristics, and player-NPC and NPC-NPC team structures.
Unlike NPCs, which are inherently embodied as avatars, co-creative tools for object building do not inherently require such embodiment. This property raises important questions about the role of embodiment in such tools and its interaction with factors like the highlighting and visualization of changes, which are the focus of this paper.
@@ 2.6. AI Contribution Visualization
Thus far, no empirical studies have investigated the effects of such an intentionally incremental visualization of AI contributions, especially not separately from the impact of AI embodiment. Therefore, we address this lack of research by contributing a study in which both embodiments and perceptual evidence in the form of incremental contribution visualization are examined in the context of a co-creative VR system.
@@ 3. Methodology
How does highlighting affect the co-creative collaboration experience, in particular regarding the predictability of AI? Based on decades of HCI research and design practices that emphasize the importance of highlighting focus or activity, we hypothesize that highlighting will positively affect perceived communication quality, as it is a well-known mechanism to guide users’ attention. For the same reason, we also expect that with highlighting, users will pay more attention to the AI, and the clarity of the AI intentions will be perceived better.
@@
How does the embodiment of an AI affect the co-creative collaboration experience, in particular regarding the perceived supportiveness, efficiency, and competence of the AI? Research on NPCs and games highlights embodiment as a critical factor in interaction, making embodiment a potentially promising factor for co-creative tools in VR. We thus expect that embodiment will affect the perceived supportiveness, efficiency, and competence of the AI, the perceived value and creativity of AI outputs, measures concerning collaborative experience, appeal, and behavioral engagement, as well as the previously mentioned contribution attribution and users’ perceived closeness to the AI.
Showing the different IVs with an example of a chair. Left are four images illustrating the difference between disembodied AI. Right are four images showing the VR scene with an embodied AI
@@ 3.1. Task
At the beginning of a modeling session, a 3D object appears on the podium, and the timer starts running. As long as the timer was running, the user could modify the object by either painting its texture (see Figure 2), adding elements to the existing object (see Figure 2), or instructing the AI to modify (see Figure 2). We chose painting as the users’ primary design responsibility as it allowed for a consistent engagement with the 3D object during the AI-driven modifications and ensured comparable levels of object alteration and creative expression while avoiding interference. For the modifications, users were always controlling if and when any changes to the object should be executed (including being able to revert the AI’s modification). We simulated the AI agent in a Wizard-of-Oz approach, in the sense that its contributions to the 3D objects were not the results of an actual generative process but instead object additions and replacements created by human 3D artists in advance.
Further, the creative contributions of the AI agent are not designed to align with the contributions of the human partner but instead emerge completely independently of the human’s behavior. This contribution style is typical for what Rezwana and Maher (Rezwana and Maher, 2022a) refer to as provoking agents (as opposed to pleasing agents), which evoke creativity in their human partners by providing divergent ideas and alternative explorations. While the AI was active, the user could not give the AI another modification instruction.